OpenStudy (jamanorthi):

Yolanda sold total 50 ticket. Adult ticket cost 5 dollar each. Student ticket cost 3 dollar each.she total collected 180 dollar. How many adult ticket did she collected? and how many student ticket did she collected?

OpenStudy (jamanorthi):

Someone please help me

OpenStudy (jamanorthi):

Yeah

OpenStudy (wolf1728):

Okay first set up some equations. 5A + 3S = $180 A + S = 50

OpenStudy (jamanorthi):

Then what I do

OpenStudy (wolf1728):

Now we have to solve these 2 equations. Multiply the bottom equation by -3 and it becomes -3A -3S = -50 We then add this to the first equation 5A + 3S = $180 -3A -3S = -150 Can you take it from there?

OpenStudy (jamanorthi):

I still don't get it How did u do it?

OpenStudy (wolf1728):

Okay the first equation sets up the money part of it $5 * (# of adult tickets) + $3*(# of student tickets) = $180

OpenStudy (wolf1728):

The second equation is concerned with the number of tickets # of Adult tickets + # of Student tickets = 50 tickets

OpenStudy (jamanorthi):

But how do I solve that problem?

OpenStudy (wolf1728):

Okay time for that. When you have 2 unknowns, you need 2 equations and you have to know how to solve those. 5A + 3S = $180 A + S = 50 One way to do this is to multiply one equation by another number so when added to the other equation one unknown will be eliminated. So we get the second equation and multiply it by -3. Why? Because this eliminates one of the unkowns. 5A + 3S = $180 -3A -3S = -150 Can you add those 2 equations?

OpenStudy (jamanorthi):

Do you have any other way to do it I am still confused with this way But thanks for your help I appreciate it If you want to help me contact me with email please My email is jamanorthi@yahoo.com

OpenStudy (wolf1728):

That's about the easiest way. Anyway, I'll send you an E-Mail.

OpenStudy (jamanorthi):

Thank you so much

OpenStudy (jamanorthi):

Can I have your email So if I need help I can ask u It's okay if you don't want to

OpenStudy (wolf1728):

I already sent you an E-Mail so just go to your mailbox.

OpenStudy (jamanorthi):

It didn't say anything about the problem

OpenStudy (wolf1728):

Well I did say a few postings ago can you add these 2 equations? 5A + 3S = $180 -3A -3S = -150

OpenStudy (jamanorthi):

I got the answer But how can I explain it in word

OpenStudy (jamanorthi):

Why did u multiply -3 with the bottom equation

OpenStudy (jamanorthi):

How did u get -3

OpenStudy (wolf1728):

It was done to eliminate 1 unknown.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (jamanorthi):

Then please answer it

OpenStudy (jamanorthi):

Someone contact me in email and help me out please